OpenStudy (anonymous):

f(x)=3x^3+4x^2-x+C and f(2)=1 what is the value of c

OpenStudy (anonymous):

put in '2' for 'x' and set the whole thing equal to '1'... then find C

OpenStudy (anonymous):

ok im trying it

OpenStudy (anonymous):

so I get fx=3(8)+4(4)-x+c

myininaya (myininaya):

Replace all x with 2 So you should have f(2)=3(8)+4(4)-2+c which can be simplified more so

myininaya (myininaya):

Then set that =1 and solve for c

OpenStudy (anonymous):

you would get : fx=3(8)+4(4)-2+c = 1 now c is easy to find,right?

OpenStudy (anonymous):

im trying to work it

OpenStudy (anonymous):

c = 1 - 24 - 16 + 2 which is: -37

OpenStudy (tasreeb):

C=-37